4. Here is a function that is either a demand function or a supply function (but not both): A change occurs so that the following function now represents the situation: We can conclude that (circle the appropriate conclusion on the answer sheet).

a. demand has increased b. demand has decreased c. supply has increased d. supply has decreased e. quantity supplied has decreased f. quantity demanded has decreased g. quantity demanded has increased h. quantity supplied has increased

6. Circle your choice for the quantity that will maximize total revenue for the function in 5 (above).

1 3 4.5 5 6 7.5

7. Suppose the price elasticity of demand for bread is 0.20. If the price of bread rises by 10%, the quantity demanded will decrease by: a. 2 percent and total expenditures on bread will rise. b. 2 percent and total expenditures on bread will fall. c. 20 percent and total expenditures on bread will rise. d. 20 percent and total expenditures on bread will fall. e. 20 percent and total expenditures on bread will be unchanged

8. Suppose that a 10% increase in income causes a 20% decrease in demand for good X. The coefficient of income elasticity of demand is: a. negative and therefore these goods are inferior goods. b. negative and therefore these goods are complements. c. positive and therefore these goods are substitutes. d. positive and therefore these goods are normal goods.

9. If price decreases from $450 to $350 and quantity demanded increases from 1,200 to 1,500, the price elasticity of demand in this price range is: 0.39 0.5 .89 1.0 1.75 2.83

10. If the elasticity of supply for crude oil is 2.5, how much will production have to decrease to match a 10% price decrease? 2.5% 8% 10% 15% 25% 40%
Answer #1

7. a. 2 percent and total expenditure on bread will rise.

8. a. negative and therefore these goods are inferior goods. This is so because the change in income decreases the quantity demanded of good X.

9. Initial price, Po= 450

Initial quantity demanded, Qo= 1200

final price, P1= 350, final quantity demanded Q1= 1500

Change in price= P1-Po= -100

Change in quantity demanded= Q1-Qo= 300

The elasticity of demand= 450/1200*300/100

= 1.125 . Thus elasticity of demand is 1.125. The options are incorrect.

10. The answer is 25%. The elasticity of supply= ratio of the percentage change in quantity supplied to change in price.

Therefore, 2.5= quantity demanded/ 10% which is equal to 25%.

Please like the answer and appreciate my effort. Please feel free to ask any doubt through the comment box. Thanks! Keep posting questions.

The other questions are not clear hence could not answer.